OpenStudy (anonymous):

Find a5 for the arithmatic sequence a1=1 and a3=25

OpenStudy (unklerhaukus):

\[a_1=1\]\[a_2=\frac{1+25}2\]\[a_3=25\] \[a_n=(a_2-a_1)n=(a_3-a_2)n\]

OpenStudy (anonymous):

whats n

OpenStudy (unklerhaukus):

\(n\) is the position of the term in the sequence \(a_1\) is the first term, \(a_2\) is the second term, \(a_5\) is the fifth term \(a_n\) is the \(n\)-th term

OpenStudy (anonymous):

so n is 5

OpenStudy (anonymous):

It will be easy if you find Arithmetic Mean first..

OpenStudy (anonymous):

\[AM = a_2 = \frac{a_1 + a_3}{2} \implies \frac{1 + 25}{2} \implies a_2 = 13\]

OpenStudy (anonymous):

Solution: a1=1 a3 = 25 a+2d = 25 2d = 24 d = 12 a5 = a + 4d = 1 + (4*12) = 49. Ans

OpenStudy (anonymous):

@curiousshubham Perhaps you shouldn't have given the direct answer to the question?

OpenStudy (anonymous):

Now you can easily find common difference here: \[d = a_2 - a_1 = 13 - 1 = 12\]

OpenStudy (anonymous):

@RolyPoly How I am supposed to provide the answer?

OpenStudy (anonymous):

And fifth you can find as: \[a_5 = a_1 + (5-1)(d)\]

OpenStudy (anonymous):

According to the Code of Conduct ( http://openstudy.com/code-of-conduct) Give Help, , Not Answers. @curiousshubham

OpenStudy (anonymous):

@RolyPoly sorry for that... a5 is the 5th term and accroding to question the value of 1st and 3rd term is 1 and 25 respectively. So all you need to do is find the term difference using the given terms i.e. 1st and 3rd. And then use the formula: nth term = a + (n-1)d.